Mal is one-and-a-half times as old as Sal, and Sal is one-and-a-half
times as old as Al. How old are Al, Sal andMal if their combined
ages total 114?

Answer by mananth(16949)

Al = x
Sal = 3x/2
Mal=(3/2)*(3x/2)=9x/4
x+3x/2 + 9x/4 = 114
multiply by 4
4x+6x+9x=456
19x=456
x=24
Al = 24
Sal = 36
Mal=54